500 alpha particles of energy 5 MeV enter ionization chamber per second the energy required for ion pair creation is 25 eV. the ionization currently produced is
1)1.6×10^-11
2)9×10^-11
1
Expert's answer
2019-04-30T09:52:22-0400

Number of pairs per second is equal


"n=\frac {500 \times 5 \times 10^6 } {25} \times 1 s =10^8 pairs/ s (1)"

Electric current


"i=10^8 pairs/ s \times1.6 \times 10^{-19} C \times 1 s = 1.6 \times 10^{-11} C"

Answer:

"1.6 \times 10^{-11} C"